# IMU: Influence-guided Machine Unlearning

### v1

@@ 2.1 Machine Unlearning
Current MU algorithms typically fall into two categories: exact unlearning methods (Ginart et al. 2019; Cao and Yang 2015; Romero, Barrio, and Belanche 2007; Karasuyama and Takeuchi 2010) that use data partitioning strategies (Bourtoule et al. 2021) or only focus on traditional machine learning models like $k$-means clustering (Ginart et al. 2019), and approximate unlearning methods (Golatkar, Achille, and Soatto 2020a; Golatkar et al. 2021; Golatkar, Achille, and Soatto 2020b; Chen et al. 2023; Heng and Soh 2023; Kumari et al. 2023; Kurmanji et al. 2023; Foster, Schoepf, and Brintrup 2024; Lyu et al. 2024; Bui et al. 2024; Ko et al. 2024; Lin et al. 2024; Spartalis et al. 2025; Alberti et al. 2025) that trade perfect deletion for computational efficiency. So far, these MU methods have demonstrated effectiveness through two key metrics: (1) degraded performance on the forgetting data, and (2) maintained utility on retain and unseen set.
Gradient ascent (GA)-based MU methods (Wu, Dobriban, and Davidson 2020; Gandikota et al. 2023; Jang et al. 2022) perform unlearning by moving model parameters away from the forgetting data through gradient ascent. While simple to implement, these methods can be sensitive to hyperparameters and may interfere with retain knowledge (Jia et al. 2023; Fan et al. 2024a). Fisher unlearning methods (Golatkar, Achille, and Soatto 2020a, b; Golatkar et al. 2021) assume that the unlearned model and the retrained model are close to each other in parameter space, and formulate Fisher Forgetting, which induces unlearning by injecting noise into the parameters proportional to their relative importance in the forgetting data compared to the retain set. Influence function-based methods estimate the impact of individual data points to perform approximate unlearning. Guo et al. (2020) propose a one‐step Newton‐update procedure based on the influence function to excise the effect of specific data points from a pretrained model. Subsequent works (Mehta et al. 2022; Liu et al. 2022) focus on ways for more efficient computation over the estimation. However, these methods make strong convexity assumptions and are still computationally expensive for large-scale models (Mehta et al. 2022). Two-stage methods (Jia et al. 2023; Fan et al. 2024b; Wu and Harandi 2024) have emerged as a practical alternative, first erasing information related to the forgetting data, then fine-tuning on retain data. These methods avoid expensive computations while achieving strong empirical results, particularly when combined with sparsity techniques.
@@ 3.2 IMU
4:  $\forall{\bm{x}}\sim{\mathcal{D}}_{f}$, estimate influence value ${\mathcal{I}}^{c}({\bm{x}},{\mathcal{D}}_{f})$ via Eq. 4: $-\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}}(\mathbf{Z_{f}})\big{)}^{\top}{\bm{H}}_{{\bm{\theta}}_{c}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}}({\bm{z}})\big{)}$.
@@
6:  Compute the influence-guided loss ${\mathcal{L}}_{\texttt{IMU}}({\mathcal{D}}_{f};{\bm{\theta}}^{(t)})$ via Eq. 5: $-\mathbb{E}_{{\bm{x}}_{i}\sim{\mathcal{D}}_{f}^{{\mathcal{I}}^{c}<0}}\left[\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\mathcal{D}}_{f})\cdot\ell\left(f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right)\right]$.
@@ 4.1 Setup
Setting Method ${\mathcal{D}}_{r}$ ${\mathcal{D}}_{f}$ $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{ACC}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{ACC}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA ()($\uparrow$)( ↑ ) $W_{\texttt{dist}}(\downarrow)$ Run time (s) ()($\downarrow$)( ↓ ) CIFAR-10 Original 99.71$\pm$0.00 099.45$\pm$0.000 094.43$\pm$0.000 00.01$\pm$0.00 Retrain 00.00$\pm$0.00 099.98$\pm$0.000 094.41$\pm$0.350 01.00$\pm$0.00 00.00$\pm$0.00 IU 00.56$\pm$0.76 084.04$\pm$9.710 079.33$\pm$8.460 00.99$\pm$0.00 04.51$\pm$1.66 055$\pm$1 GA 00.48$\pm$0.10 075.89$\pm$6.280 072.43$\pm$5.830 00.99$\pm$0.00 03.59$\pm$0.39 029$\pm$0 RL 00.71$\pm$0.42 079.01$\pm$13.25 073.33$\pm$12.25 00.99$\pm$0.00 04.04$\pm$0.61 025$\pm$5 SCAR 00.50$\pm$1.03 085.78$\pm$4.150 080.99$\pm$4.070 01.00$\pm$0.00 04.80$\pm$0.62 325$\pm$11 NPO 00.20$\pm$0.28 091.63$\pm$7.23 085.84$\pm$6.660 01.00$\pm$0.00 00.24$\pm$0.03 039$\pm$2 IMU (ours) 00.02$\pm$0.05 097.68$\pm$0.830 091.67$\pm$1.160 01.00$\pm$0.00 00.11$\pm$0.05 099$\pm$0 CIFAR-100 Original 99.89$\pm$0.19 099.99$\pm$0.02 076.63$\pm$5.47 00.01$\pm$0.00 Retrain 00.00$\pm$0.00 099.53$\pm$2.79 074.05$\pm$8.95 01.00$\pm$0.00 00.00$\pm$0.00 IU 00.00$\pm$0.00 069.29$\pm$4.00 061.85$\pm$5.18 01.00$\pm$0.00 04.82$\pm$2.00 19$\pm$2 SSD 00.00$\pm$0.00 077.37$\pm$9.93 046.85$\pm$2.90 01.00$\pm$0.00 03.15$\pm$1.27 10$\pm$0 GA 00.00$\pm$0.00 068.29$\pm$5.27 040.89$\pm$8.61 01.00$\pm$0.00 05.22$\pm$0.14 09$\pm$0 RL 00.00$\pm$0.00 073.87$\pm$8.81 042.80$\pm$7.68 01.00$\pm$0.00 07.73$\pm$0.81 22$\pm$3 SCAR 00.00$\pm$0.00 077.79$\pm$1.28 060.00$\pm$1.48 01.00$\pm$0.00 07.13$\pm$1.51 68$\pm$0 NPO 00.00$\pm$0.00 090.47$\pm$3.15 055.80$\pm$7.55 01.00$\pm$0.00 05.35$\pm$1.10 25$\pm$3 IMU (ours) 00.00$\pm$0.00 098.06$\pm$1.35 063.75$\pm$7.85 01.00$\pm$0.00 02.09$\pm$0.37 14$\pm$0 CIFAR-100 Original 100.00$\pm$0.00 100.00$\pm$0.000 077.80$\pm$2.98 00.00$\pm$0.00 Retrain 00.00$\pm$0.000 099.33$\pm$1.160 076.49$\pm$2.53 01.00$\pm$0.00 00.00$\pm$0.00 IU 00.00$\pm$0.000 060.58$\pm$4.890 054.50$\pm$5.31 01.00$\pm$0.00 05.80$\pm$0.08 18$\pm$0 SSD 00.00$\pm$0.000 083.07$\pm$1.130 052.20$\pm$1.32 01.00$\pm$0.00 04.81$\pm$0.50 10$\pm$0 GA 00.00$\pm$0.000 075.56$\pm$9.460 049.50$\pm$6.16 01.00$\pm$0.00 07.54$\pm$1.29 09$\pm$0 RL 00.00$\pm$0.000 076.49$\pm$11.51 047.69$\pm$4.54 01.00$\pm$0.00 07.57$\pm$0.07 19$\pm$2 SCAR 00.00$\pm$0.000 078.72$\pm$4.800 056.55$\pm$5.31 01.00$\pm$0.00 05.88$\pm$0.25 67$\pm$0 NPO 00.00$\pm$0.000 090.22$\pm$4.570 058.90$\pm$6.34 01.00$\pm$0.00 05.61$\pm$0.25 26$\pm$2 IMU (ours) 00.00$\pm$0.000 098.61$\pm$0.850 067.60$\pm$3.44 01.00$\pm$0.00 03.69$\pm$0.16 14$\pm$0
@@ 4.2 Results on class-wise unlearning
We first evaluate on CIFAR-10, trying to forget an entire class category. As shown in Table 1, our proposed method IMU achieves superior unlearning performance. IU fails to effectively preserve the model utility, showing significant drops of $\sim$15in accuracy in both ${\mathcal{D}}_{r}$ and the test set ${\mathcal{D}}_{t}$. GA and RL perform even worse in this regard. The SOTA retain-data-free MU method SCAR demonstrates strong forgetting ability; however, it comes at the cost of notable degradation in model utility, with accuracies on ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{t}$ reduced by approximately 14%. We further employ the potent LLM unlearning method, NPO, for image classification. It significantly outperforms baselines in both the forget quality and model utility preservation.
Similarly, our proposed method IMU, not only achieves strong forget capabilities, i.e., 0.02% on ${\mathcal{D}}_{f}$, but also best preserves the model utility, with only about a 2% drop on both ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{t}$, indicating that IMU strikes a desirable balance between effective forgetting and minimal impact on model utility. In addition, IMU demonstrates favorable efficiency, is over 3$\times$ faster than SCAR.
@@ 4.3 Results on subclass-wise unlearning
[…]
@@ 4.4 Results on person re-identification
[…]
@@ 4.6 Ablation study
$r$ $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA ()($\uparrow$)( ↑ ) Run time ()($\downarrow$)( ↓ ) 1.00 0.02$\pm$0.03 97.93$\pm$0.84 91.22$\pm$0.77 1.00$\pm$0.00 70$\pm$0 0.80 0.18$\pm$0.32 97.89$\pm$0.80 91.19$\pm$0.58 1.00$\pm$0.00 63$\pm$0 0.60 0.04$\pm$0.08 97.94$\pm$0.86 91.32$\pm$0.62 1.00$\pm$0.00 62$\pm$0 0.40 0.03$\pm$0.05 98.00$\pm$0.71 91.14$\pm$0.83 1.00$\pm$0.00 60$\pm$0 0.20 0.00$\pm$0.00 98.62$\pm$0.70 91.98$\pm$0.53 1.00$\pm$0.00 58$\pm$0 0.05 0.00$\pm$0.00 99.79$\pm$0.19 93.78$\pm$0.24 1.00$\pm$0.00 57$\pm$0
@@
$\nu$ $\texttt{Acc}_{{\mathcal{D}}_{f}}(\downarrow)$ $\texttt{Acc}_{{\mathcal{D}}_{r}}(\uparrow)$ $\texttt{Acc}_{{\mathcal{D}}_{t}}(\uparrow)$ MIA ()($\uparrow$)( ↑ ) Run time ()($\downarrow$)( ↓ ) 0 0.02$\pm$0.02 97.75$\pm$0.70 91.86$\pm$0.53 1.00$\pm$0.00 076$\pm$0 111 0.00$\pm$0.00 97.65$\pm$0.73 91.71$\pm$0.54 1.00$\pm$0.00 268$\pm$1 222 0.00$\pm$0.00 97.72$\pm$0.73 91.82$\pm$0.58 1.00$\pm$0.00 175$\pm$3
[…]
@@ 6.1 Implementation details
All the experiments were performed on NVIDIA RTX3090 with Intel Xeon processors. For image classification scenario, the original model and retrained model were both trained with over 182182182 epochs using the SGD optimizer with a cosine-scheduled, and learning rate was initialized at 0.10.10.1. For both NPO and RL, training spans 101010 epochs within the interval $[10^{-4},10^{-1}]$, and $\beta$ was selected from the range [0.1,5][0.1,5][ 0.1 , 5 ]. GA’s training settings around a 555-epoch learning rate search within the interval $[10^{-5},10^{-2}]$. In the case of IU, we explore the parameter $\alpha$ within the range [1,20][1,20][ 1 , 20 ]. For SSD, searching for the parameter selection weighting $\alpha$ and dampening constant $\gamma$ is executed within the range [1,100][1,100][ 1 , 100 ] and [0.1,1][0.1,1][ 0.1 , 1 ] respectively, while searching for the learning rate within the range $[10^{-4},10^{-2}]$. For the SCAR method, the value of temperature was defined within the range [2,5][2,5][ 2 , 5 ], the parameter $\lambda_{2}$ was set as 0.010.010.01, and $\lambda_{1}$ was in the range of [1,10][1,10][ 1 , 10 ] with the learning rate searching in the interval $[10^{-6},10^{-2}]$. Learning rates was selected in the range $[10^{-4},10^{-2}]$ for SalUn and let sparsity ratios equal to 0.50.50.5. Lastly, for IMU method, we trained for 3 epochs, searching for learning rates in the range $[10^{-5},10^{-2}]$ and $l_{1}$ regularization intensity in the range $[2\times 10^{-4},2\times 10^{-2}]$. For person re-identification unlearning task, original model was trained over 606060 epochs with learning rate as $3\times 10^{-4}$. In the case of IMU, it is trained for $1.5\times 10^{-3}$ iterations with a learning rate of $1.5\times 10^{-3}$, with $\alpha$ set to 0.020.020.02. For sequence modeling unlearning problem, the original model and retrained model were both trained with over 555 epochs using softmax activation, and the learning rate was initialized at $5\times 10^{-4}$. As for SimNPO and NPO method, the value of $\beta$ is searched in interval of [0.1,5][0.1,5][ 0.1 , 5 ] with learning rate in range $[2\times 10^{-4},8\times 10^{-3}]$, while the learning rate of GA was selected from $[2\times 10^{-4},10^{-2}]$. For our IMU method, learning rate was selected as $5\times 10^{-4}$ with $\alpha$ set as 0.0020.0020.002. Additionally, all the unlearning methods were conducted over 505050 iterations. Lastly, in LLM unlearning scenario, $\beta$ was selected from [1,5][1,5][ 1 , 5 ] for NPO and SimNPO, and the learning rate was searched in range $[10^{-6},10^{-4}]$. Additionally, all methods in this task were executed over 6 epochs and distributed across 2 NVIDIA RTX 3090 GPUs. Our source code will be publicly available to support the reproducibility of the results.

### v2

@@ 2.1 Machine Unlearning
Current MU algorithms typically fall into two categories: exact unlearning methods (Ginart et al. 2019; Cao and Yang 2015; Romero, Barrio, and Belanche 2007; Karasuyama and Takeuchi 2010) that use data partitioning strategies (Bourtoule et al. 2021) or only focus on traditional machine learning models like $k$-means clustering (Ginart et al. 2019), and approximate unlearning methods (Golatkar, Achille, and Soatto 2020a; Chen et al. 2023; Heng and Soh 2023; Kumari et al. 2023; Kurmanji et al. 2023; Foster, Schoepf, and Brintrup 2024; Lyu et al. 2024; Bui et al. 2024; Ko et al. 2024; Lin et al. 2024; Spartalis et al. 2025; Alberti et al. 2025; Wu et al. 2025; Wu and Harandi 2024a) that trade perfect deletion for computational efficiency. So far, these MU methods have demonstrated effectiveness through two key metrics: (1) degraded performance on the forgetting data, and (2) maintained utility on retain and unseen set.
Gradient ascent (GA)-based MU methods (Wu, Dobriban, and Davidson 2020; Gandikota et al. 2023; Jang et al. 2022) perform unlearning by moving model parameters away from the forgetting data through gradient ascent. While simple to implement, these methods can be sensitive to hyperparameters and may interfere with retain knowledge (Jia et al. 2023; Fan et al. 2024a). Fisher unlearning methods (Golatkar, Achille, and Soatto 2020a, b; Golatkar et al. 2021) assume that the unlearned model and the retrained model are close to each other in parameter space, and formulate Fisher Forgetting, which induces unlearning by injecting noise into the parameters proportional to their relative importance in the forgetting data compared to the retain set. Influence function-based methods estimate the impact of individual data points to perform approximate unlearning. Guo et al. (2020) propose a one‐step Newton‐update procedure based on the influence function to excise the effect of specific data points from a pretrained model. Subsequent works (Mehta et al. 2022; Liu et al. 2022) focus on ways for more efficient computation over the estimation. However, these methods make strong convexity assumptions and are still computationally expensive for large-scale models (Mehta et al. 2022). Two-stage methods (Jia et al. 2023; Fan et al. 2024b; Wu and Harandi 2024b) have emerged as a practical alternative, first erasing information related to the forgetting data, then fine-tuning on retain data. These methods avoid expensive computations while achieving strong empirical results, particularly when combined with sparsity techniques.
@@ 3.2 IMU
4:  $\forall{\bm{x}}\sim{\mathcal{D}}_{f}$, estimate influence value ${\mathcal{I}}^{c}({\bm{x}},{\mathcal{D}}_{f})$ via Eq. 4: $\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}}(\mathbf{Z_{f}})\big{)}^{\top}{\bm{H}}_{{\bm{\theta}}_{c}}^{-1}\nabla_{{\bm{\theta}}_{c}}\ell\big{(}h_{{\bm{\theta}}_{c}}({\bm{z}})\big{)}$.
@@
6:  Compute the influence-guided loss ${\mathcal{L}}_{\texttt{IMU}}({\mathcal{D}}_{f};{\bm{\theta}}^{(t)})$ via Eq. 5: $\mathbb{E}_{{\bm{x}}_{i}\sim{\mathcal{D}}_{f}^{{\mathcal{I}}^{c}<0}}\left[\hat{{\mathcal{I}}^{c}}({\bm{x}}_{i},{\mathcal{D}}_{f})\cdot\ell\left(f_{{\bm{\theta}}}\left({\bm{x}}_{i}\right)\right)\right]$.
@@ 4.2 Results on class-wise unlearning
We first evaluate on CIFAR-10, trying to forget an entire class category. As shown in Table 1, our proposed method IMU achieves superior unlearning performance. IU fails to effectively preserve the model utility, showing significant drops of $\sim$15% in accuracy in both ${\mathcal{D}}_{r}$ and the test set ${\mathcal{D}}_{t}$. GA and RL perform even worse in this regard. The SOTA retain-data-free MU method SCAR demonstrates strong forgetting ability; however, it comes at the cost of notable degradation in model utility, with accuracies on ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{t}$ reduced by approximately 14%. We further employ the potent LLM unlearning method, NPO, for image classification. It significantly outperforms baselines in both the forget quality and model utility preservation.
[…]
@@ 4.3 Results on subclass-wise unlearning
Despite the increased difficulty, IMU consistently demonstrates better performance than other MU methods. As shown in Table 1, all MU methods completely remove the knowledge about ${\mathcal{D}}_{f}$, while IMU best preserving the performance over ${\mathcal{D}}_{r}$, having an accuracy around $98\%$ and achieves a high test accuracy compared to other baselines.
@@ 4.4 Results on person re-identification
Pedestrians’ body shapes, clothing, facial features, and other biometric characteristics are often sensitive and uniquely identifiable, making them important targets for protection in MU scenarios. To this end, we also evaluate MU methods on the person re-identification task. This task involves matching images of the same individual captured under varying camera viewpoints or multiple cameras. Here, we aim to forget all samples associated with a particular identity (e.g., $\texttt{pid}=1$). This task poses unique challenges to MU methods, as forgetting a specific individual requires precise removal of identity-related features while preserving general person-level recognition performance.
We adopt GradCAM (Selvaraju et al. 2017) to visualize regions where models focus on w/o and w/ IMU. As shown in Figure 3, when evaluated on ${\mathcal{D}}_{r}$, IMU preserves focus on discriminative yet identity-agnostic regions. In contrast, when evaluated on ${\mathcal{D}}_{f}$, our scrubbed model significantly shifts attention away from identity-revealing regions (e.g., clothing logos, face, and hairstyle), indicating successful removal of sensitive cues. In Table 2, we further evaluate model utility by measuring mAP and Top-$k$ accuracy on a query-gallery split consisting of individuals entirely unseen during training. In general, IMU achieves a good trade-off between forget quality and model utility.
@@ 4.5 Case study: Markov chains
Aside from vision tasks, we further evaluate our method on a simple sequence modeling problem. Following the setting in SimNPO (Fan et al. 2024a), we construct a mixture of Markov chains with a state space of size 10, the retain distribution consists of Markov chains that transition uniformly among states $\{1,2,3\}$, while the forget distribution is a mixture of two Markov chains Forget1 (transition uniformly among $\{4,5,6\}$) and Forget2 (transition uniformly among $\{7,8,9\}$) with equal probability. Each sequence has length $T$ and is denoted as $s=(s_{1},s_{2},\cdots,s_{T})$ where $s_{t}\in\{0,1,\cdots,9\}$ for $t\in[0,T]$. A GPT-2 model is trained to learn the conditional distribution. Details can be found in §7 of SimNPO (Fan et al. 2024a). Table 3 presents results compared to NPO and SimNPO. Performance is evaluated by the loss values over the retain and forget sets, as well as the respective KL divergence distance calculated compared with the retrained model. Our method surpasses NPO and SimNPO in both forget quality and model utility.
4.6 Ablation study
@@
[…]
@@
We finally conduct ablation studies to analyze the effects of two key design choices in our method: (1) the ratio $r$ of top-ranked forgetting data selected based on influence values, and (2) the update frequency $v$ of influence scores during the unlearning process. Table 4 presents the effect of varying the top-$r$ fraction of the forgetting data ${\mathcal{D}}_{f}$ for unlearning. As $r$ decreases, the model retains strong forget quality as evidenced by near-zero accuracy on ${\mathcal{D}}_{f}$, while achieving slightly improved accuracy on ${\mathcal{D}}_{r}$ and ${\mathcal{D}}_{t}$. Notably, using only 5% of the most influential data points is sufficient to induce forgetting, while also enhancing generalization and utility, highlighting the efficiency of targeting high-impact data points. Table 5 investigates the impact of varying the frequency $v$ of influence score updates. All variants achieve strong forget quality, with minor differences in model utility. Infrequent updates are sufficient for maintaining effective unlearning, suggesting that IMU is robust.
@@ 6.1 Implementation details
All the experiments were performed on NVIDIA RTX3090 with Intel Xeon processors. For image classification scenario, the original model and retrained model were both trained with over $182$ epochs using the SGD optimizer with a cosine-scheduled, and learning rate was initialized at $0.1$. For both NPO and RL, training spans $10$ epochs within the interval $10^{-4},10^{-1}]$, and $\beta$ was selected from the range $[0.1,5]$. GA’s training settings around a $5$-epoch learning rate search within the interval $10^{-5},10^{-2}]$. In the case of IU, we explored the parameter $\alpha$ within the range $[1,20]$. For SSD, searching for the parameter selection weighting $\alpha$ and dampening constant $\gamma$ was executed within the range $[1,100]$ and $[0.1,1]$, respectively, while searching for the learning rate within the range $10^{-4},10^{-2}]$. For the SCAR method, the value of temperature was defined within the range $[2,5]$, the parameter $\lambda_{2}$ was set as $0.01$, and $\lambda_{1}$ was in the range of $[1,10]$ with the learning rate searching in the interval $10^{-6},10^{-2}]$. Learning rate was selected in the range $10^{-4},10^{-2}]$ for SalUn and let sparsity ratio equal to $0.5$. Lastly, for IMU method, we trained for 3 epochs, searching for learning rates in the range $10^{-5},10^{-2}]$ and $l_{1}$ regularization intensity in the range $2\times 10^{-4},2\times 10^{-2}]$. For person re-identification unlearning task, original model was trained over $60$ epochs with learning rate as $3\times 10^{-4}$. In the case of IMU, it is trained for $1.5\times 10^{-3}$ iterations with a learning rate of $1.5\times 10^{-3}$, with $\alpha$ set to $0.02$. For sequence modeling unlearning problem, the original model and retrained model were both trained with over $5$ epochs using softmax activation, and the learning rate was initialized at $5\times 10^{-4}$. As for SimNPO and NPO method, the value of $\beta$ was searched in interval of $[0.1,5]$ with learning rate in range $2\times 10^{-4},8\times 10^{-3}]$, while the learning rate of GA was selected from $2\times 10^{-4},10^{-2}]$. For our IMU method, learning rate was selected as $5\times 10^{-4}$ with $\alpha$ set as $0.002$. Additionally, all the unlearning methods were conducted over $50$ iterations. Lastly, in LLM unlearning scenario, $\beta$ was selected from $[1,5]$ for NPO and SimNPO, and the learning rate was searched in range $10^{-6},10^{-4}]$. Additionally, all methods in this task were executed over 10 epochs and distributed across 2 NVIDIA RTX 3090 GPUs. The official code is available at: https://github.com/goodluckisallyouneed/IMU.git.